# A Graph-Based Laser Path Solver Algorithm for Virtual Reality Laboratory Simulations

### v1

@@ 4.2 Experiment-specific performance analysis
Although femtoPro’s open-world approach allows for a wide range of custom experimental configurations, we included various predefined scenarios, or setups, designed for educational purposes that represent typical use cases aligned with standard curricula of established textbooks [11, 23, 24]. The layout of every experiment is tailored to specific educational objectives, spanning from basic (e.g., adjusting optical elements) to more advanced tasks (e.g., measuring nonlinear autocorrelation, see Fig. 2).
@@
Given that the memory usage remains well within available capacity (see Section 4.1) and it is unlikely that any experimental setup will approach this limit, the following analysis will primarily focus on time efficiency of CPU operations. For virtual reality experiences, maintaining a minimum frame rate of 120 fps is crucial for minimizing nausea and discomfort [14]. However, since most currently available mobile VR headsets, including those targeted by femtoPro, only support up to 90 fps, we establish this as our minimum acceptable frame rate, corresponding to a time budget of ca. 11 ms per frame ($\frac{1000~{}\text{ms}}{90}=11~{}\text{ms}$).
For mobile platforms, game engines such as Unity advise reserving 35 % or more of the available frame time for CPU idling. This approach aids in controlling device heat and preserving battery life [25]. At 90 fps, this equates to approximately 4 ms of idle time per frame. Based on performance profiling on our target platform, we additionally set aside a sufficiently large buffer of 3 ms for system components, such as sensory input processing. As a result, the maximum time budget per frame that femtoPro can allocate to the simulation code is approximately 4 ms.
@@ 4.3 Considerations on user experience
Motion sickness in VR is primarily caused by discrepancies between physical sensations of movement and visual input lagging behind due to a low and/or unstable framerate [14]. The susceptibility and severity of motion sickness is increased when the head is moved rapidly, such as during movement through the VR environment. Contrary to this, during simulation interaction where the head remains in a relatively fixed position with the eyes focusing on specific elements, the impact of an unstable framerate is less pronounced. Selective updating is beneficial in this context, as it conserves potentially scarce system resources especially when the user is most susceptible to motion sickness.
@@
Lower CPU resource usage, resulting from minimal simulation workload also reduces battery drain on mobile VR devices. The effect becomes increasingly pronounced with the proportion of time the simulation operates under a reduced workload. The impact of selective updating on power consumption was assessed using the Meta Quest mobile VR headset, which represents the lower end of supported devices, featuring a battery capacity of 14 Wh [26, 27]. With the simulation using all available resources each frame, power draw continuously reached 14 W. Under the assumption that CPU resources are not internally throttled at lower battery levels to conserve power, this results in a total playtime of approximately one hour from a fully charged to a completely drained battery state. Conversely, with the simulation continuously in idle mode, the power draw was decreased by ~15 %, resulting in 10 more minutes of playtime before the battery is depleted. Consequently, by implementing the selective updating strategy, the frequency and length of recharging breaks is reduced, ensuring longer play sessions without interruption, and thus, offering a smoother experience for the user.
@@ 5 Conclusion and future work
We implemented a custom-tailored path solver algorithm to meet femtoPro’s requirements and associated challenges. Every frame, the graph’s topology is updated to reflect any changes introduced by user interactions. This is achieved through a node processing loop that successively generates each node’s outgoing edges by applying linear and nonlinear optical transformations to its incident edges. With femtoPro’s open-world approach, beams can be aligned in various ways that might, in an extreme case, create infinite cycles of reflecting laser light, presenting a significant challenge to real-time calculations. To address this challenge, we implemented a cap on the number of optical responses that can be processed per frame.
@@ 5.3 Conflict of interest
“femtoPro” is a registered trademark of the University of Würzburg. The software is available for free for both educational and commercial usage via the femtoPro website [20]. There are no commercial interests or partnerships established.
@@ 5.6 Data availability
The raw data that support the findings of this paper are openly available in Zenodo at https://doi.org/10.5281/zenodo.15533468 [28].
@@ 5.7 Materials availability
A demonstration trailer of the femtoPro app is available at the femtoPro website [20].
@@ 5.8 Code availability
The latest femtoPro version 0.7.0 features 26 tutorial levels and is freely available for download on the femtoPro website [20].

### v2

@@ 4.2 Experiment-specific performance analysis
Although femtoPro’s freely configurable and scalable lab environment allows for a wide range of custom experimental configurations, we included various predefined scenarios, or setups, designed for educational purposes that represent typical use cases aligned with standard curricula of established textbooks[trager2012, diels2006, wolfgang2013]. The layout of every experiment is tailored to specific educational objectives, spanning from basic (e.g., adjusting optical elements) to more advanced tasks (e.g., measuring nonlinear autocorrelation, see Fig. 3).
@@
Given that the memory usage remains well within available capacity (see Section 4.1) and it is unlikely that any experimental setup will approach this limit, the following analysis will primarily focus on time efficiency of CPU operations. For virtual reality experiences, maintaining a minimum frame rate of 120 fps is crucial for minimizing nausea and discomfort [effectOfFramerate2023]. However, since most currently available mobile VR headsets, including those targeted by femtoPro, only support up to 90 fps, we establish this as our minimum acceptable frame rate, corresponding to a time budget of ca. 11 ms per frame ($\frac{1000~\text{ms}}{90}=11~\text{ms}$).
For mobile platforms, game engines such as Unity advise reserving 35 % or more of the available frame time for CPU idling. This approach aids in controlling device heat and preserving battery life [frameIdleBudget2024]. At 90 fps, this equates to approximately 4 ms of idle time per frame. Based on performance profiling on our target platform, we additionally set aside a sufficiently large buffer of 3 ms for system components, such as sensory input processing. As a result, the maximum time budget per frame that femtoPro can allocate to the simulation code is approximately 4 ms.
@@ 4.3 Considerations on user experience
Motion sickness in VR is primarily caused by discrepancies between physical sensations of movement and visual input lagging behind due to a low and/or unstable framerate [effectOfFramerate2023]. The susceptibility and severity of motion sickness is increased when the head is moved rapidly, such as during movement through the VR environment. Contrary to this, during simulation interaction where the head remains in a relatively fixed position with the eyes focusing on specific elements, the impact of an unstable framerate is less pronounced. Selective updating is beneficial in this context, as it conserves potentially scarce system resources especially when the user is most susceptible to motion sickness.
@@
Lower CPU resource usage, resulting from minimal simulation workload also reduces battery drain on mobile VR devices. The effect becomes increasingly pronounced with the proportion of time the simulation operates under a reduced workload. The impact of selective updating on power consumption was assessed using the Pico 3 mobile VR headset featuring a battery capacity of 19.61 Wh [pico3Specs]. With the simulation using all available resources each frame, power draw continuously reached 11.7 W. Under the assumption that CPU resources are not internally throttled at lower battery levels to conserve power, this results in a total playtime of approximately 1 h 41 min from a fully charged to a completely drained battery state. Conversely, with the simulation continuously in idle mode, the power draw was decreased by ca. 10 %, resulting in 11 more minutes of playtime before the battery is depleted. Consequently, by implementing the selective updating strategy, the frequency and length of recharging breaks is reduced, ensuring longer play sessions without interruption, and thus, offering a smoother experience for the user.
@@ 5 Conclusion and future work
We implemented a custom-tailored path solver algorithm to meet femtoPro’s requirements and associated challenges. Every frame, the graph’s topology is updated to reflect any changes introduced by user interactions. This is achieved through a node processing loop that successively generates each node’s outgoing edges by applying linear and nonlinear optical transformations to its incident edges. With femtoPro’s freely configurable lab environment, beams can be aligned in various ways that might, in an extreme case, create infinite cycles of reflecting laser light, presenting a significant challenge to real-time calculations. To address this challenge, we implemented a cap on the number of optical responses that can be processed per frame.
@@ 5.3 Conflict of interest
“femtoPro” is a registered trademark of the University of Würzburg. The software is available for free for both educational and commercial usage via the femtoPro website [femtoProWebsite]. There are no commercial interests or partnerships established.
@@ 5.6 Data availability
The raw data that support the findings of this paper are openly available in Zenodo at https://doi.org/10.5281/zenodo.15533468 [zenodo2025].
@@ 5.7 Materials availability
A demonstration trailer of the femtoPro app is available at the femtoPro website [femtoProWebsite].
@@ 5.8 Code availability
The latest femtoPro version 0.7.0 features 26 tutorial levels and is freely available for download on the femtoPro website [femtoProWebsite].